OpenStudy (anonymous):

An unbiased coin is tossed. If the result is a head, a pair of unbiased dice is rolled and the sum of the numbers obtained is noted. If the result is a tail, a well shuffled card of eleven cards numbered 2,3,4...12 is picked and the number on the card is noted. The probability that the noted number is either 7 or 8 is...

OpenStudy (anonymous):

is the question considering the "noted number" to be either the sum of the dice roll OR the card chosen?

OpenStudy (anonymous):

there are 2 cases out of 11 for the card choice, which only occurs 1/2 of the time, i.e. when tails is the coin flip. But there are a bunch of 7 and 8 options for a 2-dice roll. Interesting question :)

OpenStudy (anonymous):

It can be both, the sum when its head and the card when its tail..

OpenStudy (anonymous):

there are 6 ways to roll a 7, and 5 ways to roll an 8. So out of 36 outcomes, 11 are "winners". I think you have (1/2)(11/36) for the dice option and (1/2)(2/11) for the cards option.

OpenStudy (anonymous):

Yes..that's right for both the parts...Now the final answer? @JakeV8

OpenStudy (anonymous):

it's the sum of the two outcomes, since either produces a winner. P = 11/72 + 2/22 = 242/1584 + 144/1584 = 386/1584 = 0.2437

OpenStudy (anonymous):

Is this your work or just Saturday edition of "fun with probability"? :)

OpenStudy (anonymous):

Won't it be the multiplication of the two probabilities....? :)

OpenStudy (anonymous):

I was thinking that through too :) I might have messed that up.

OpenStudy (zarkon):

your answer is correct jake

OpenStudy (anonymous):

Yes you are right..thanks..Its not work. I like probability. So, I solve them regularly.

OpenStudy (anonymous):

Since either path produces winners with the given probabilities, then they add.

OpenStudy (anonymous):

That's my "final answer", Regis.

OpenStudy (anonymous):

No its not multiplication...its addition

OpenStudy (anonymous):

It was great practice for me :) Thanks for posting it :)

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

Ohh thanks .. I too have some probability questions .. Hope to post it soon..

OpenStudy (anonymous):

Shoot them sooner..lets try @Miyuru

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I have to drop off in a minute.. will be sorry to miss them :) Hey @Miyuru, maybe for fun (loosely defined), you can throw an @JakeV8 in the questions... I'd like to read them later.

OpenStudy (anonymous):

OK sure @JakeV8 I would mention u on my question hope to see u soon..bye

OpenStudy (anonymous):

bye all, thanks again...

OpenStudy (anonymous):

U are welcome @JakeV8 U did a great job for us...